OpenStudy (anonymous):

What is .336 written as a fraction

OpenStudy (anonymous):

Write the decimal number as a fraction (over 1) 0.336 = 0.336 / 1 Multiplying by 1 to eliminate 3 decimal places we multiply top and bottom by 3 10's Numerator (N) N = 0.336 × 10 × 10 × 10 = 336 Denominator (D) D = 1 × 10 × 10 × 10 = 1000 N / D = 336 / 1000 Simplifying our fraction = 336/1000 = 42/125

OpenStudy (anonymous):

boxes are multiply signs